# Safe In-Context Reinforcement Learning

### v1

@@ Abstract
In-context reinforcement learning (ICRL) is an emerging RL paradigm where the agent, after some pretraining procedure, is able to adapt to out-of-distribution test tasks without any parameter updates. The agent achieves this by continually expanding the input (i.e., the context) to its policy neural networks. For example, the input could be all the history experience that the agent has access to until the current time step. The agent’s performance improves as the input grows, without any parameter updates. In this work, we propose the first method that promotes the safety of ICRL’s adaptation process in the framework of constrained Markov Decision Processes. In other words, during the parameter-update-free adaptation process, the agent not only maximizes the reward but also minimizes an additional cost function. We also demonstrate that our agent actively reacts to the threshold (i.e., budget) of the cost tolerance. With a higher cost budget, the agent behaves more aggressively, and with a lower cost budget, the agent behaves more conservatively.
@@ 1 Introduction
Reinforcement learning (RL, Sutton & Barto (2018)) is a framework for solving sequential decision making problems via trial and error. Modern RL agents are usually parameterized by deep neural networks, and the trial-and-error process is typically achieved by updating the parameters of the neural networks (Mnih et al., 2015). Recently, in-context reinforcement learning (ICRL, Moeini et al. (2025)) emerges as an RL paradigm that achieves this trial-and-error process without any neural network parameter updates (Duan et al., 2016; Mishra et al., 2018; Xu et al., 2022; Kirsch et al., 2023; Lin et al., 2023; Lu et al., 2023; Raparthy et al., 2023; Sinii et al., 2023; Grigsby et al., 2024a; b; Cook et al., 2024; Dai et al., 2024; Dong et al., 2024; Elawady et al., 2024; Huang et al., 2024; Krishnamurthy et al., 2024; Laskin et al., 2023; Lee et al., 2023; Shi et al., 2024; Wang et al., 2025a; Zisman et al., 2023; 2024; Song et al., 2025; Wang et al., 2025b)
Specifically, the pretrained agent neural network takes as input not only the observation at the current time step but also an additional context. For example, the simplest choice of the context can be all the history of experiences that the agent has access to up to the current time step (Laskin et al., 2023). It is then observed that the agent’s performance improves as the length of the context grows, even if the neural network parameters are kept fixed and even if the agent is evaluated on a task that it has never encountered during the pretraining process (Laskin et al., 2023). This phenomenon is called ICRL. A widely accepted hypothesis for this performance improvement is that the forward pass of the pretrained agent network implicitly implements some RL algorithm during inference time to process the information in the context (Laskin et al., 2023; Lin et al., 2023; Wang et al., 2025a; b). As the context length grows, the amount of information grows, and the implemented RL algorithm in the forward pass is then able to output better actions.
ICRL has recently received increasing attention given its cheap yet strong adaptation capability. The adaptation is cheap in that it does not require parameter tuning. So to support such adaptation, the infrastructure only needs to support the inference of the neural network. For example, by using ICRL in LLMs’ inference time, the LLMs are able to significantly improve on challenging tasks such as ScienceWorld (Wang et al., 2022), HMMT (Harvard–MIT Mathematics Tournament, ), and AIME (Mathematical Association of America, ) without any fine-tuning (Song et al., 2025). The adaptation is strong in that ICRL allows adaptation to not only unseen but also out-of-distribution tasks. For example, Laskin et al. (2023) show that after pretraining on some bandit tasks, their ICRL agent is able to adapt quickly to test bandit tasks that have opposite optimal arms to the pretraining bandit tasks. Kirsch et al. (2023) show that after pretraining on Ant robot manipulation tasks, their ICRL agent is able to adapt to pole balancing tasks.
Despite the empirical success and the promising potential, the safety of ICRL has long been ignored. Here we discuss safety in the framework of constrained Markov Decision Processes (Altman, 2021), where each action the agent takes not only incurs a reward but also incurs a cost. Then when an agent adapts to a new task, it needs to not only maximize the rewards but also minimize the cost, all without updating the parameters.
We argue that safe ICRL is a necessary condition to enable many possible real-world applications of ICRL, such as embodied AI. Although such safe decision-making problems have been well studied in the classical RL setup (Garcıa & Fernández, 2015), to our knowledge, no prior work has studied the safety of ICRL, which is the gap that this paper shall close. We envision that a safe ICRL agent should have the following two desiderata, both of which should be achieved without any parameter updates during the adaptation to new tasks.
  • When evaluated in out-of-distribution tasks, the agent should be able to maximize the rewards while keeping the cost below a user-specified threshold.
  • The agent should actively react to the cost threshold. For example, a smaller cost threshold should induce a more conservative behavior, while a larger cost threshold should induce a more aggressive behavior.
@@
The key contribution of this work is to design the first safe ICRL agent that simultaneously fulfills the two desiderata. Specifically, we make the following two contributions.
  • We study both off-line (cf. (Laskin et al., 2023)) and on-line pretraining approaches (cf. (Duan et al., 2016; Wang et al., 2016; Grigsby et al., 2024a; b)) for safe ICRL. We identify key limitations of the offline approach and propose a novel online pretraining approach that fulfills the two desiderata.
  • We empirically validate our proposed approach in challenging safe decision making problems. Particularly, our test task is not only unseen during the pretraining stage but also out of the pretraining task distribution. As a result, to succeed in our benchmarks, the agent has to demonstrate strong out-of-distribution generalization. Particularly, the agent has to explore efficiently in out-of-distribution test tasks, all without parameter updates.
@@ 2 Background
We consider a finite horizon Markov Decision Process (MDP, Bellman (1957); Puterman (2014)) with a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a transition function $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$, an initial distribution $p_{0}:\mathcal{S}\to[0,1)$, and a horizon length $T$. At time step 0, an initial state $S_{0}$ is sampled from $p_{0}$. At time step $t$, an agent at a state $S_{t}$ selects an action $A_{t}$ according to its policy $\pi:\mathcal{A}\times\mathcal{S}\to[0,1]$, i.e., $A_{t}\sim\pi(\cdot|S_{t})$. The agent then proceeeds to a successor state $S_{t+1}\sim p(\cdot|S_{t},A_{t})$ and obtains a reward $R_{t+1}\doteq r(S_{t},A_{t})$. This process continues until the time T1T-1, where the agent executes the action $A_{T-1}$ and receives the last reward $R_{T}$. We use $\tau\doteq(S_{0},A_{0},R_{1},S_{1},\dots,S_{T-1},A_{T-1},R_{T})$ to denote an episode. We use $k$ to index episodes and $t$ to index time steps within an episode. When necessary, we write $S_{t}^{k}$ to denote the state at time step $t$ within the $k$-th episode. $A_{t}^{k}$ and $R_{t}^{k}$ are similarly defined. For any episode $\tau$, we use $G(\tau)\doteq\sum_{t=1}^{T}R_{t}$ to denote the return associated with this episode. The performance of a policy $\pi$ is then measured by the expected total rewards $J(\pi)\doteq\mathbb{E}_{\tau\sim\pi}[G(\tau)]$. One fundamental task in RL, called control, is to adapt the policy $\pi$ to maximize $J(\pi)$.
Modern RL methods solve the control problem via parameterizing the policy $\pi$ with neural networks (Mnih et al., 2015). We denote such parameterized policy as $\pi_{\theta}$, where $\theta$ denotes the parameters of the neural network. Typically, $\theta$ is updated at every time step, e.g., $\theta_{t+1}$ is obtained by updating $\theta_{t}$ with the newly obtained information $(S_{t},A_{t},S_{t+1},R_{t+1})$. The performance of the policy $\pi_{\theta_{t}}$ improves along with such parameter updates. In other words, the reinforcement learning process occurs with the parameter updates. ICRL is an emerging paradigm that allows the reinforcement learning process to occur without any parameter updates. In ICRL, the policy $\pi$ takes an additional context as input. We denote the context at time $t$ within the $k$-th episode as $H_{t}^{k}$ and express the policy as $\pi(\cdot|S_{t}^{k},H_{t}^{k})$ to emphasize such dependency. The construction of the context is an active research area and we refer the reader to Moeini et al. (2025) for a detailed survey. The simplest example is to use all the history as context, e.g., $H_{t}^{k}\doteq(\tau_{1},\dots,\tau_{k-1},S_{0}^{k},A_{0}^{k},R_{1}^{k},S_{1}^{k},\dots S_{t-1}^{k},A_{t-1}^{k},R_{t}^{k})$. After some pretraining procedure on multiple MDPs to be detailed soon, we can obtain a parameter $\theta_{*}$. It is observed that the performance of $\pi_{\theta_{*}}(\cdot|S_{t}^{k},H_{t}^{k})$ improves as the context grows with $\theta_{*}$ kept fixed. In other words, the reinforcement learning process now occurs without any parameter updates. This cannot be attributed to the hypothesis that $\theta_{*}$ memorizes a good policy, as such performance improvement is also observed even when $\pi_{\theta_{*}}$ is evaluated in out-of-distribution MDPs that are very different from the pretraining MDPs (Laskin et al., 2023; Kirsch et al., 2023). A widely accepted hypothesis is that the neural network $\theta_{*}$ implicitly implements some RL algorithm in its forward pass, such that it can process the context with the RL algorithm in the inference time (Moeini et al., 2025). As the context grows, the inference-time RL algorithm in the forward pass gains access to more information. So the performance of the policy improves. This hypothesis is also theoretically backed by Lin et al. (2023); Wang et al. (2025a; b).
@@
We now elaborate on the pretraining methods for ICRL. According to the taxonomy in Moeini et al. (2025), the pretraining can be divided into supervsied pretraining (Laskin et al., 2023; Lin et al., 2023) and reinforcement pretraining (Duan et al., 2016; Wang et al., 2016; Grigsby et al., 2024a; b; Wang et al., 2025a; b). Supervised pretraining is essentially behavior cloning but the goal is to imitate an algorithm instead of a policy. Supervised pretraining is usually done in an offline manner. Specifically, by running an existing RL algorithm on an MDP, we can collect a sequence of episodes, denoted as $\Xi\doteq(\tau_{1},\tau_{2},\dots,\tau_{K})$. We call $\Xi$ a trajectory. Suppose the RL algorithm behaves well, we would expect that the episode return $G(\tau_{k})$ increases as $k$ increases. So the trajectory $\Xi$ demonstrates the learning process of the RL algorithm in the MDP. In supervised pretraining, we collect multiple trajectories by running multiple existing RL algorithms on multiple MDPs, yielding a dataset $\quantity{\Xi_{i}}$. The policy $\pi_{\theta}$ is then trained with an imitation learning loss. In other words, for a state $S_{t}^{k}$ from a trajectory $\Xi_{i}$ in the dataset, the loss for updating $\theta$ is
@@
By asking the neural network to imitate the behavior of RL algorithms demonstrated in the dataset, we expect the neural network to be able to distill some RL algorithm into its forward pass. This is known as algorithm distillation (Laskin et al., 2023). Instead of distilling existing RL algorithms, reinforcement pretraining asks the network to discover its own RL algorithm. Reinforcement pretraining is typically done in an online manner. At time step $t$ within the $k$-th episode, the loss is
@@
3 Safe In-Context Reinforcement Learning
While ICRL has demonstrated remarkable generalization in out-of-distribution tasks, the safety of such generalization has been largely overlooked. This paper is the first to study the safety of ICRL. Particularly, we use the constrained MDP (CMDP) framework (Altman, 2021). In addition to the reward function $r$, we also have a cost function $c:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. At a state $S_{t}$, the agent executes an action $A_{t}$ and receives both a reward $R_{t+1}$ and a cost $C_{t+1}\doteq c(S_{t},A_{t})$. An episode is then
@@ 3 Safe In-Context Reinforcement Learning
We use $G_{c}(\tau)\doteq\sum_{t=1}^{T}C_{t}$ to denote the toal cost in the episode $\tau$. After some pretraining procedure, we obtain a policy $\pi_{\theta_{*}}$. We then execute this policy in a test MDP for multiple episodes $\tau_{1},\dots,\tau_{K}$. With ICRL, we can expect that the episode return $G(\tau_{k})$ increaes as $k$ increaes. But for safe ICRL, we would like to additionally see that the episode total cost $G_{c}(\tau_{k})$ decreases as $k$ increases, ideally below some user-given threshold, say $\delta$. Moreover, when the user is more tolerant of the cost (i.e., with a larger $\delta$), the agent should obtain higher rewards (i.e., larger $G(\tau_{k})$). We formalize these desiderata as the following constrained problem:
$\textstyle\max_{\theta}\mathbb{E}_{\pi_{\theta}}[\sum_{k=1}^{K}G(\tau_{k})]\mbox{\quad s.t.\quad}\forall k,\mathbb{E}_{\pi_{\theta}}\quantity[G_{c}(\tau_{k})]\leq\delta.$ (4)
Here, $K$ is the number of episodes that are allowed in the test time. $\tau_{k}$ is the $k$-th epsiode sampled from the policy $\pi_{\theta}$. The policy $\pi_{\theta}$ can depends on the context (i.e., $\pi_{\theta}(A_{t}^{k}|S_{t}^{k},H_{t}^{k})$) but the neural network parameters $\theta$ is fixed for all the $K$ episodes. We note that (4) is defined on a single test MDP that the algorithm has no access to during pretraining. This test MDP is usually sufficiently different from the MDPs that the algorithm has access to during pretraining. As a result, the agent cannot solve this test MDP via memorizing some good policies in the pretraining MDPs. Instead, it has to perform reinforcement learning over the $K$ episodes, although the parameters $\theta$ are fixed. Notably, although we formulate our problem in MDPs, the tasks we work on are only partially observable. This is a widely used practice in the RL community to promote the clarity of presentation (Mnih et al., 2015). We now explore different pretraining methods for solving the above problem.
Safe Supervised Pretraining.
We first establish safe supervised pretraining as a baseline approach. Similar to naive supervised pretraining, we collect a dataset $\quantity{\Xi_{i}}$ by executing various safe RL algorithms (e.g., Tessler et al. (2018); Ray et al. (2019)) in various safety-sensitive tasks (e.g., Ji et al. (2023); Gu et al. (2025)). Now each trajectory $\Xi_{i}$ consists of multiple episodes and each episode contains both the reward and the cost (cf. (3)). To further guide the action generation, we additionally condition the policy on both the return-to-go (RTG) $G_{t}(\tau)$ and the cost-to-go (CTG) $G_{c,t}(\tau)$. RTG is the total future rewards in the current episode, i.e., $G_{t}(\tau)\doteq\sum_{i=t+1}^{T}R_{t}$, and CTG is the total future cost in the current episode, i.e., $G_{c,t}(\tau)\doteq\sum_{i=t+1}^{T}C_{t}$. The RTG and CTG are inspired by Liu et al. (2023) for constrained decision Transformers and are designed to bias the agent towards achieving the specified return and cost. The main difference from Liu et al. (2023) is that in addition to RTG and CTG, we use entire episodes in input, while Liu et al. (2023) only includes states and actions but do not include per-step reward and per-step cost. As a result, Liu et al. (2023) still misses the opportunity to learn the algorithm demonstrated in the dataset, and they demonstrated only limited generalization capability. Specifically, given a trajectory $\Xi=(\tau_{1},\dots,\tau_{K})$ in the dataset, the safe supervised pretraining loss at a state $S_{t}^{k}$ is
$-\log\pi_{\theta}(A_{t}^{k}|S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k})).$ (5)
We recall that the context $H_{t}^{k}$ is defined as $(\tau_{1},\dots,\tau_{k-1},S_{0}^{k},A_{0}^{k},R_{1}^{k},C_{1}^{k},\dots,S_{t-1}^{k},A_{t-1}^{k},R_{t}^{k},C_{t}^{k})$. In the pretraining, we do have access to RTG and CTG since the episode is already complete. But in the test time, we do not have access to them since they are calculated based on future rewards and costs, which are not available yet. Instead, we replace them with a spectrum of target RTG and CTG values. This gives the user the opportunity to control the trade-off between rewards and costs in the test time without fine-tuning the parameters. For example, in the test time, if the ratio RTG / CTG is small, it means the user is more risk-tolerant. If the ratio RTG / CTG is large (in the extreme case, CTG = 0), it means the user is more safety-focused.
Although the proposed safe supervised pretraining approach is to our knowledge novel, it is indeed a straightforward extension of existing unconstrained superivsed pretraining approaches like Laskin et al. (2023). We regard this as a baseline approach for sanity check. Any meaningful safe ICRL method should at least outperform this baseline.
Safe Reinforcement Pretraining.
Our safe supervised learning follows the offline pretraining paradigm. It is well-known that the performance of offline trained policies is usually worse than the online trained ones when online training is plausible (Vinyals et al., 2019; Mathieu et al., 2021). Motivated by this, we now study safe reinforcement pretraining. We note that even when we consider online pretraining, we still do not have access to the test MDP where (4) is defined. By online pretraining, we mean that we can interact with the pretraining MDPs directly. By contrast, offline supervised pretraining has access to only datasets collected from pretraining MDPs.
The underlying hypothesis of reinforcement pretraining is that by minimizing the loss (2) in a wide range of pretraining MDPs, the sequence model should be able to discover its own RL algorithm and implement it in the forward pass (see Lu et al. (2023); Grigsby et al. (2024a; b) for empirical evidence and Wang et al. (2025b) for theoretical evidence). Following this hypothesis, we now design an algorithm to solve (4) on a single pretraining MDP. To solve (4), primal-dual approaches in safe-RL (Achiam et al., 2017; Tessler et al., 2018; Liang et al., 2018; Ray et al., 2019) convert it to the dual problem below [^1] We omit $\theta$ and use $\max_{\pi}f(\pi)$ instead of $\max_{\theta}f(\pi_{\theta})$ for simplicity.:
$\textstyle\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})=\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}\quantity[{\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}-\sum_{k=1}^{K}\lambda_{k}({\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]}-\delta)]$ (6)
where ${\bm{\lambda}}\succeq 0$ is a componentwise inequality for the Lagrangian multipliers. Notably, the function class of the policy $\pi$ in $\max_{\pi}$ is all the policies in the form of (5) but without RTG, i.e., the policy can depend on the context and the CTG.
However, we observe that while the CMDP formulation requires a separate constraint and Lagrangian multiplier for each evaluation episode $\tau_{k}$, this is not desirable in our problem, as it would need to fix the number of evaluation episodes at the beginning of pretraining. Using fewer episodes would result in some multipliers being updated fewer times than others, which makes the optimization unsable and using more episodes is not possible.
Furthermore, there should be symmetry among all the constraints since they are all coming from a single cost function but simply correspond to different episodes. These encourage the use of a single multiplier for all episodes. However, when using a single multiplier, we would need a new way to not over-penalize the policy in the episodes that are already satisfying the constraint. To this end, We propose a modified Lagrangian function and an iterative optimization scheme and show it is more stable empirically while satisfying the desirable optimization properties. Let $g_{k}(\pi)\coloneqq{\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]}-\delta$. We consider the following surrogate objective for the policy
$\textstyle L_{\Sigma}(\pi,\lambda)={\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}-\lambda\sum_{k=1}^{K}[g_{k}(\pi)]_{+}$

### v2

@@ Abstract
In-context reinforcement learning (ICRL) is an emerging RL paradigm where an agent, after pretraining, can adapt to out-of-distribution test tasks without any parameter updates, instead relying on an expanding context of interaction history. While ICRL has shown impressive generalization, safety during this adaptation process remains unexplored, limiting its applicability in real-world deployments where test-time behavior is expected to be safe. In this work, we propose SCARED: Safe Contextual Adaptive Reinforcement via Exact-penalty Dual, the first method that promotes safe adaptation of ICRL under the constrained Markov decision process framework. During the parameter-update-free adaptation process, our agent not only maximizes the reward but also keeps the accumulated cost within a user-specified safety budget. We also demonstrate that the agent actively reacts to the safety budget; with a higher safety budget, the agent behaves more aggressively, and with a lower safety budget the agent behaves more conservatively. Across challenging benchmarks, SCARED consistently enables safe and robust in-context adaptation, outperforming existing ICRL and safe meta-RL baselines.
In-context Reinforcement Learning, Safe Reinforcement Learning
@@ 1 Introduction
Reinforcement learning (RL, Sutton & Barto (2018)) has achieved remarkable success in sequential decision-making, from game playing (Mnih et al., 2015) to robotic control (Levine et al., 2016). These achievements rely on deep neural networks whose parameters are updated through trial-and-error interaction with the environment. However, parameter updates at test time are not always feasible for adaptation: gradient computation may be unavailable, careful hyperparameter calibration can be required, and fine-tuning in meta-learning settings risks catastrophic forgetting of previously learned behaviors (Kirkpatrick et al., 2017).
Recently, in-context reinforcement learning (ICRL, Moeini et al. (2025)) has emerged as an RL paradigm that achieves this trial-and-error process without any neural network parameter updates (Duan et al., 2016; Xu et al., 2022; Lee et al., 2023; Laskin et al., 2023; Raparthy et al., 2023; Sinii et al., 2023; Huang et al., 2024; Krishnamurthy et al., 2024). Specifically, ICRL enables agents to adapt to new tasks at test time by conditioning on their interaction history, without updating network parameters. This allows a pretrained model to generalize to unseen environments with improving performance as more contextual experience accumulates.
Using ICRL in real-world deployments is especially desirable because, to enable adaptation, the infrastructure only needs to support the inference of the neural network. Despite the empirical success and the promising potential of ICRL (Laskin et al., 2023; Kirsch et al., 2023; Dai et al., 2024; Dong et al., 2024), the safety of ICRL has long been overlooked, which is a necessary condition for many possible real-world applications, such as embodied AI. In standard RL, agents are deployed on tasks they are trained for and expected to satisfy the safety constraints enforced during training. In contrast, in ICRL, agents explore and learn new tasks during test time. Therefore, safety constraints must be satisfied not just by the final learned behavior, but also during the adaptation process (Xu & Zhu, 2025). To our knowledge, no prior work has studied the safety of ICRL, which is the gap that this paper aims to close.
We model the safe ICRL framework as a constrained Markov decision process (CMDP, Altman (2021)), where agents receive not only a reward but also a cost signal after each action. We envision that safe ICRL agents should maximize the rewards while keeping the cost below a user-specified cost budget even when evaluated in out-of-distribution (OOD) tasks. In addition, they should actively react to the cost budget; e.g., a smaller cost budget should induce more conservative behavior for safety, while a larger cost budget should allow more aggressive behavior for reward maximization. Meeting these desiderata is challenging; existing ICRL methods have no mechanism to incorporate cost signals, while safe meta-RL methods use past trial history only to fine-tune parameters, which does not capture the precise details of past experiences.
@@
The key contribution of this work is to design the first safe ICRL agent that simultaneously fulfills the aforementioned desiderata, all without any parameter updates during the adaptation to new tasks. Specifically, we make the following contributions:
  • Safe ICRL Problem Formulation. We formalize safe ICRL under the CMDP framework, where agents adapt to new tasks at test time without parameter updates while adjusting to safety budgets.
  • OOD Benchmarks for Safe In-Context Adaptation. We introduce benchmark environments that evaluate safe exploration and extrapolative generalization beyond interpolation regimes under safety constraints.
  • SCARED: Safe Contextual Adaptive Reinforcement via Exact-Penalty Dual. We propose SCARED, an online reinforcement pretraining method that satisfies CMDP safety constraints. We show SCARED outperforms the baselines, increasing cumulative reward and decreasing cost as context grows, while adapting to the safety budgets.
@@ 2.1 Markov Decision Processes
We model the interaction with an environment for a given task as a finite-horizon MDP (Puterman, 2014). An MDP is defined by a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, a transition function $p:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$, an initial distribution $p_{0}:\mathcal{S}\to[0,1]$, and a horizon length $T$. An agent interacts with the environment by following a policy $\pi:\mathcal{A}\times\mathcal{S}\to[0,1]$. Specifically, after starting from an initial state $S_{0}\sim p_{0}$, at any time step $t$, the agent takes an action $A_{t}\sim\pi(\cdot|S_{t})$, observes a reward $R_{t+1}\doteq r(S_{t},A_{t})$, and proceeds to a next state $S_{t+1}\sim p(\cdot|S_{t},A_{t})$. This process continues until the time T1T-1, when the agent takes a final action $A_{T-1}$ and receives the last reward $R_{T}$. We use $\tau\doteq(S_{0},A_{0},R_{1},S_{1},\dots,S_{T-1},A_{T-1},R_{T})$ to denote an episode, and use $k$ to index episodes and $t$ to index time steps in an episode. We write $S_{t}^{k}$, $A_{t}^{k}$, and $R_{t}^{k}$ to denote the state, action, and reward at time step $t$ in the $k$-th episode, respectively. For any episode $\tau$, we define its return as $G(\tau)\doteq\sum_{t=1}^{T}R_{t}$. The performance of a policy $\pi$ is then measured by the expected total rewards $J(\pi)\doteq\mathbb{E}_{\tau\sim\pi}[G(\tau)]$ obtained under $\pi$. The fundamental task in MDPs is to obtain an optimal policy $\pi$ that maximizes $J(\pi)$, which typically requires a learning-based approach when a transition model is absent.
2.2 Reinforcement Learning
The goal of RL is to find an optimal policy for a given task by learning through interaction with the environment (Sutton & Barto, 2018). Classic approaches often construct tabular value functions that are used to explicitly map each state to an action, therefore are feasible only when the state and action spaces are discrete and small. Recent RL techniques, combined with modern deep learning methods can handle large or continuous state–action spaces via function approximation with neural networks parameterized by weights $\theta$ (Wang et al., 2022). Starting from an initial policy $\pi_{\theta_{0}}$, deep RL methods typically update the parameters using observed transitions $(S_{t},A_{t},S_{t+1},R_{t+1})$ to obtain $\pi_{\theta_{t+1}}$. Performance can improve through these updates, which aim to reduce prediction errors via backpropagation through a loss function. These techniques have been successful at learning high-performing policies for a particular task; however, their performance can degrade substantially across OOD tasks and environments. In such settings, policies are often re-learned or fine-tuned, which limits the large-scale deployment of a single trained policy.
2.3 In-Context Reinforcement Learning
Recently, ICRL has emerged as a paradigm that allows RL agents to adapt to new tasks at test time without requiring any parameter updates after pretraining. In ICRL, the parameters $\theta_{*}$ are first learned with a pretraining procedure on multiple MDPs. The resulting policy $\pi_{\theta_{*}}$ parameterized by $\theta_{*}$ is then deployed to adapt to new tasks by taking context as additional input. We denote the context at time $t$ within the $k$-th episode as $H_{t}^{k}$ and we write the policy as $\pi(\cdot|S_{t}^{k},H_{t}^{k})$ to make this dependence explicit. A simple choice is to use the history of transitions as context (see Moeini et al. (2025) for alternative approaches); i.e., $H_{t}^{k}\doteq(\tau_{1},\dots,\tau_{k-1},\tau_{k}^{t})$ where $\tau_{k}^{t}\doteq(S_{0}^{k},A_{0}^{k},R_{1}^{k},S_{1}^{k},\dots S_{t-1}^{k},A_{t-1}^{k},R_{t}^{k})$ is the observed prefix of the current episode. Adaptation to a new task in ICRL occurs as the performance of $\pi_{\theta_{*}}(\cdot|S_{t}^{k},H_{t}^{k})$ improves with growing context while $\theta_{*}$ remains fixed. This generalization capability is hypothesized to be a consequence of the neural network implicitly implementing an RL algorithm in its forward pass, processing the context at inference time (Laskin et al., 2023; Kirsch et al., 2023). As the context grows, this inference-time algorithm gains access to more information, improving policy performance. This hypothesis is also theoretically supported by Lin et al. (2023); Wang et al. (2025a, b). Despite the remarkable generalization of ICRL to OOD tasks, the safety of such generalization has been overlooked.
@@ 2.3 In-Context Reinforcement Learning
We now discuss the pretraining methods for ICRL, which can typically be divided into supervised pretraining (Laskin et al., 2023; Lin et al., 2023) and reinforcement pretraining (Duan et al., 2016; Wang et al., 2016; Grigsby et al., 2024a, b; Wang et al., 2025a, b). Supervised pretraining employs behavior cloning with the goal of imitating an algorithm rather than a policy, and it is usually done in an offline manner. Specifically, by running an existing RL algorithm on an MDP, we can collect a sequence of episodes, denoted as $\Xi\doteq(\tau_{1},\tau_{2},\dots,\tau_{K})$. We call $\Xi$ a trajectory. The episode return $G(\tau_{k})$ is expected to increase as $k$ increases; therefore, the trajectory $\Xi$ can be used to demonstrate the learning progression of the RL algorithm on the MDP. In supervised pretraining, multiple trajectories are collected by running multiple existing RL algorithms on multiple MDPs, yielding a dataset $\quantity{\Xi_{i}}$. The policy $\pi_{\theta}$ is then trained with an imitation learning loss. In other words, for a state $S_{t}^{k}$ from a trajectory $\Xi_{i}$ in the dataset, the loss for updating $\theta$ is
@@
By optimizing $\theta$ to imitate the behavior of RL algorithms demonstrated in the dataset, the policy $\pi_{\theta}$ is expected to embed some RL-like mechanism into its forward pass, which is known as algorithm distillation (Laskin et al., 2023). Instead of distilling existing RL algorithms, reinforcement pretraining optimizes the network to maximize the return on many different MDPs, guiding it to discover its own RL algorithm by providing the interaction history. Reinforcement pretraining is typically done in an online manner. At time step $t$ within the $k$-th episode, the loss is
@@
3 Problem Formulation: Safe ICRL
We study the safety of ICRL using the CMDP framework (Altman, 2021). In addition to the reward function $r$, a CMDP involves a cost function $c:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ with an associated user-given budget $\delta$. At each time step $t$, after taking an action $A_{t}$ at a state $S_{t}$, the agent additionally receives a cost $C_{t+1}\doteq c(S_{t},A_{t})$. An episode in a CMDP therefore becomes
@@ 3 Problem Formulation: Safe ICRL
We denote the total cost of an episode $\tau$ by $G_{c}(\tau)\doteq\sum_{t=1}^{T}C_{t}$, which should, in expectation, remain below the budget $\delta$.
The objective of ICRL is to pretrain a set of parameters $\theta_{*}$ such that the resulting policy $\pi_{\theta_{*}}$ maximizes the sum of expected returns for a new task at test time. This is usually achieved by the return $G(\tau_{k})$ of an episode $\tau_{k}$ being increased with $k$, as the policy learns more about the environment. In safe ICRL, we additionally expect the episode cost $G_{c}(\tau_{k})$ to decrease as $k$ increases, ideally remaining below a user-specified cost budget $\delta$.
We formalize this safe ICRL objective as follows:
$\textstyle\max_{\theta}\mathbb{E}_{\pi_{\theta}}[\sum_{k=1}^{K}G(\tau_{k})]\text{~~s.t.~~}\forall k,\mathbb{E}_{\pi_{\theta}}\quantity[G_{c}(\tau_{k})]\leq\delta.$ (4)
Here, $K$ is the number of total test episodes, and $\tau_{k}$ denotes the $k$-th test episode obtained by executing the policy $\pi_{\theta}$ on a novel CMDP that is not accessible during pretraining. Throughout test time, the CMDP and the policy parameters $\theta$ are fixed; therefore, the policy $\pi_{\theta}(\cdot|S_{t}^{k},H_{t}^{k})$ can adapt to the CMDP only by utilizing the context, i.e., the history of test transitions. Although we formulate our problem using CMDPs, our formulation can be used for partially observable systems without loss of generality.
4 SCARED: Safe Contextual Adaptive Reinforcement Pretraining
We now introduce SCARED, our online reinforcement pretraining approach for the safe ICRL problem. SCARED enables safe in-context adaptation to unseen CMDPs by regulating episode-level costs within a given budget. We establish a pretraining procedure that minimizes the loss (2) across a range of CMDPs, extending the generalization capability from reward maximization to also respecting the safety budget, solely based on context without parameter updates at test time. Specifically, we design an algorithm to solve (4) by converting the primal problem to its dual form [^1] We omit $\theta$ and write $\max_{\pi}f(\pi)$ instead of $\max_{\theta}f(\pi_{\theta})$ for brevity.:
$\hskip-5.0pt\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})=$
$\hskip-5.0pt\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}\quantity[{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]}{-}\sum_{k=1}^{K}\hskip-1.0pt\lambda_{k}({\mathbb{E}_{\pi}\hskip-1.0pt\quantity[G_{c}(\tau_{k})]}{-}\delta)].$ (5)
where ${\bm{\lambda}}\succeq 0$ denotes componentwise nonnegativity of the Lagrangian multipliers.
The maximization over $\pi$ ranges over contextual policies that condition not only on the interaction history but also on the total cost expected over the remainder of the episode. We refer to this total future cost as cost-to-go (CTG) and define it as $G_{c,t}(\tau)\doteq\sum_{i=t+1}^{T}C_{i}$. At the start of each episode, CTG is set to the budget $\delta$, i.e., $G_{c,0}=\delta$. Policies then sample actions conditioned on CTG as $A_{t}^{k}\sim\pi_{\theta}\!\left(\cdot\mid S_{t}^{k},H_{t}^{k},G_{c,t}(\tau_{k})\right)$.
In safe ICRL, applying a separate constraint and Lagrange multiplier to each evaluation episode $\tau_{k}$ is undesirable as this requires fixing the number of episodes at the start of pretraining and leads to uneven update frequencies across multipliers. For example, if we expect $K$ evaluation episodes, we must initialize at least $K$ multipliers. If pretraining rollouts contain fewer episodes than the number of multipliers, later multipliers are updated less frequently, which destabilizes the optimization (see Figure 5.(c)). Furthermore, the constraints should be symmetric since they are associated with a single cost function but simply correspond to different episodes. These motivate the use of a single multiplier for all episodes. However, using a single multiplier will over-penalize the policy on episodes that already satisfy the constraint. To prevent this, we propose a modified Lagrangian function and an iterative optimization scheme. We show this approach is empirically more stable while preserving the desirable optimization properties.
Let $g_{k}(\pi)\coloneqq{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[G_{c}(\tau_{k})]}-\delta$. We consider the following surrogate objective for the policy
$\textstyle L_{\Sigma}(\pi,\lambda)={\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]}-\lambda\sum_{k=1}^{K}[g_{k}(\pi)]_{+}$